# Gate teleportation-assisted routing for quantum algorithms

### v1

@@ 4 Conclusions
Our evaluation of routing performance on benchmark circuits—including the Deutsch-Jozsa algorithm, graph state, QFT, and QAOA—demonstrated that teleportation-assisted routing can reduce circuit depth by $10$ to $20\%$ compared to a native line topology. While teleporting CNOT gates was generally effective, our case study on QAOA circuits showed that teleporting RZZ gates using controlled-U teleportation provided better performance. These findings highlight the potential of teleported gates to improve routing efficiency and emphasize the importance of customizing the routing strategy to align with specific gate requirements and device topology. Despite these promising results, certain limitations were identified. The overhead introduced by auxiliary qubits during teleportation must be carefully managed to ensure the practicality of this method for current quantum devices. While the total number of CNOT gates in the circuit is relatively high, the number of CNOTs acting on the data qubits remains comparable to that of a line topology. This is particularly relevant since the auxiliary qubits are solely used for teleportation and are discarded after mid-circuit measurements. Additionally, incorporating error mitigation techniques for these measurements could further enhance performance [39], making teleportation-based routing more viable on noisy intermediate-scale quantum (NISQ) devices.
@@ 5 Acknowledgements
We are grateful for useful discussion with Nicola Lo Gullo and the whole TORQS project members. The work is supported by Business Finland through the project ”Towards Reliable Quantum Software Development: Approaches and Use Case” (TORQS), Grants No. 8582/31/2022 and 8436/31/2022. M.S. acknowledges H2Future project through the Research Council of Finland (Grant. No. 352788) and the University of Oulu, and the Research Council of Finland Grant No. 316619.

### v2

@@ 4 Conclusions
Our evaluation of routing performance on benchmark circuits—including the Deutsch-Jozsa algorithm, graph state, QFT, and QAOA—demonstrated that teleportation-assisted routing can reduce circuit depth by $10$ to $20\%$ compared to a native line topology. While teleporting CNOT gates was generally effective, our case study on QAOA circuits showed that teleporting RZZ gates using controlled-U teleportation provided better performance. These findings highlight the potential of teleported gates to improve routing efficiency and emphasize the importance of customizing the routing strategy to align with specific gate requirements and device topology. Despite these promising results, certain limitations were identified. The overhead introduced by auxiliary qubits during teleportation must be carefully managed to ensure the practicality of this method for current quantum devices. While the total number of CNOT gates in the circuit is relatively high, the number of CNOTs acting on the data qubits remains comparable to that of a line topology. This is particularly relevant since the auxiliary qubits are solely used for teleportation and are discarded after mid-circuit measurements. The scaling of auxiliary qubit usage depends on algorithmic requirements and the number of available qubits on the hardware, making it an adaptable strategy rather than a fixed overhead. Additionally, incorporating error mitigation techniques for these measurements could further enhance performance [45], making teleportation-based routing more viable on noisy intermediate-scale quantum (NISQ) devices. Note that the choice of the initial layout could also affect the results, and the RTG method could utilize existing routing and mapping techniques to optimize the initial layout based on the given algorithm and hardware topology. In a more advanced routing scheme, potential teleportation paths could also be considered during initial layout selection to further enhance efficiency. Alternatively, a dedicated subset of qubits within the hardware could be specifically assigned for teleportation, ensuring that idle qubits are optimally positioned to facilitate nonlocal operations.
@@ 5 Acknowledgements
We are grateful for useful discussion with Nicola Lo Gullo and the whole TORQS project members. The work is supported by Business Finland through the project ”Towards Reliable Quantum Software Development: Approaches and Use Case” (TORQS), Grants No. 8582/31/2022 and 8436/31/2022,, and the project ”Securing the Quantum Software Stack” (SeQuSoS), Grants No. 112/31/2024. M.S. acknowledges H2Future project through the Research Council of Finland (Grant. No. 352788) and the University of Oulu, and the Research Council of Finland Grant No. 316619.